At the central Perk coffee shop, Destiny sold 30 more cups of cappuccino than expresso, for a total of 178.50 worth of expresso and cappuccino. How many cups of each were sold if expresso cost 2.00 per cup and cappuccino cost 2.50 per cup?
Answer by richwmiller(17219)

2e+2.5c=178.5
e+30=c
2e+2.5(e+30)=178.5
2e+2.5e+2.5*30=178.5
4.5e+75=178.5
4.5e=103.5
e=103.5/4.5
e=23
c=53
check
46+2.5*53=178.5
2.5*53=132.5
106+26.5=132.5
ok